\Mr. R. Plane, a retired college administrator, consumes only grapes and the composite good C (Pc = $1). His income consists of $10,000/yr from social security, plus the proceeds from whatever he sells of the 2000 bushels of grapes he harvests annually from his vineyard. Last year, grapes sold for $2/bushel, and Mr. Plane consumed all 2000 bushels of his grapes (did not sell any) in addition to 10,000 units of C. This year, the price of grapes is $3/bushel, while Pc remains $1. If his indifference curves have the conventional shape, will this year’s consumption of grapes be greater than, smaller than, or the same as last year’s? Will this year’s consumption of C be greater than, smaller than, or the same as last year’s? Explain briefly and graph.
Answer #1

Consider the given problem here there are two goods “G=grapes” and “C=composite good”. Now, given the information provided the initial budget line is “Pc*C + Pg*G = 10,000 + Pg*2000”, => “C + 2*G = 10,000 + 2*2000”.

=> C + 2*G = 14,000.

In the above fig the initial budget line is “A1B1” and the initial equilibrium is “E1(G=2000, C=10,000)”. Now, as the price of grapes increases to “3” the new budget line is “C+3*G = 16,000”. In the above fig the new budget line is “A2B2” which is much steeper and passes through “E1”, => the optimum consumption point “E2” lies on higher indifference curve.

So, as the price of grapes increases “Mr. Plane” become better off.